# Multi-state Models For Modeling Disease Histories Based On Longitudinal Data

### v1

@@ Abstract
Multi-stage disease histories derived from longitudinal data are becoming increasingly available as registry data and biobanks expand. Multi-state models are suitable to investigate transitions between different disease stages in presence of competing risks. In this context, however their estimation is complicated by dependent left-truncation, multiple time scales, index event bias, and interval-censoring. In this work, we investigate the extension of piecewise exponential additive models (PAMs) to this setting and their applicability given the above challenges. In simulation studies we show that PAMs can handle dependent left-truncation and accommodate multiple time scales. Compared to a stratified single time scale model, a multiple time scales model is found to be less robust to the data generating process. We also quantify the extent of index event bias in multiple settings, demonstrating its dependence on the completeness of covariate adjustment. In general, PAMs recover baseline and fixed effects well in most settings, except for baseline hazards in interval-censored data. Finally, we apply our framework to estimate multi-state transition hazards and probabilities of chronic kidney disease (CKD) onset and progression in a UK Biobank dataset (n=$142,667$). We observe CKD progression risk to be highest for individuals with early CKD onset and to further increase over age. In addition, the well-known genetic variant rs77924615 in the UMOD locus is found to be associated with CKD onset hazards, but not with risk of further CKD progression.
@@ 1 Introduction
As the adoption of registry data expands [1, 2] and biobanks gain importance [3, 4, 5], multi-stage disease histories are increasingly becoming available for use in both clinical routine and epidemiological research. Examples include Chronic Kidney Disease (CKD), where disease stages are defined by thresholds of serum creatinine-based estimated glomerular filtration rate (eGFR) measurements [6]; diabetes, defined based on fasting plasma glucose levels [7]; or age-related macular degeneration (AMD), characterized by the accumulation of subretinal drusenoid deposits (early/intermediate AMD) and cell atrophy (late AMD).
In this context, it is usually of interest to view transitions and risk factors conditional on current disease stage: For example, given onset of Mild CKD at 50 years of age and progression to Severe CKD 10 years later, what is the probability of developing end-stage kidney disease (ESKD)? And is a genetic variant a risk factor for CKD onset, for CKD progression, or for both? It is appealing to employ multi-state models to derive transition probabilities and risk factor associations of multi-stage disease histories, such as CKD stages (Figure 1).
@@
Multi-stage disease histories can be formalized via state diagrams, defined by all possible transitions $o\rightarrow\ell$ (where $o$ and $\ell$ are the from- and to-state, respectively), indexed by $k=1,\ldots,q$. In multi-state modeling, hazard rates $h(t\mid\mathbf{x}_{i}):=\lim_{\Delta t\to 0}\frac{1}{\Delta t}P(t\leq T<t+\Delta t\mid T\geq t,\mathbf{x}_{i})$ are extended to transition-specific hazards
$h_{k}(t\mid\mathbf{x}_{i,k}):=\lim_{\Delta t\to 0}\frac{1}{\Delta t}P(t\leq T<t+\Delta t,o\mid T\geq t,\ell,\mathbf{x}_{i,k}),$ (1)
where $i\in\{1,\ldots,n\}$ is the subject index and $\mathbf{x}_{i,k}$ the subject- and transition-specific covariate vector.
Right-censored, single-event survival data typically comes in tuples $(y_{i},d_{i},\mathbf{x}_{i})$, where $y_{i}$ is the observed event or censoring time and $d_{i}$ is the status indicator. Multi-state data encodes transitions via tuples $(y^{entry}_{i,k},y^{exit}_{i,k},d_{i,k},\mathbf{x}_{i,k})$, where $y^{entry}_{i,k}$ and $y^{exit}_{i,k}$ are risk set entry and exit times, due to either transitioning or censoring, and $d_{i,k}$ is the transition-specific status indicator. Such multi-state data enables the computation of the transition probability matrix [8, 9], whose $(\ell,o)$-th element contains the probability of transition $k$, as matrix product across previous transition probability matrices. Since transition probabilities can be time-inhomogeneous, we define the instantaneous probability $dP^{k}(t)=(\textbf{dP}(t))_{\ell,o}=P(t\leq T\leq t+\Delta t,o\mid T\geq t,\ell),\ell\neq o$. The probability of staying in state $\ell$ is $(\textbf{dP}(t))_{\ell,\ell}=1-\sum_{k}dP^{k}(t)$ and the transition probability matrix can be defined as
@@
where u∈[s,t) denotes the product integral over the interval $[s,t)$. In practice, $dP^{k}(t)$ is approximated by $H_{k}(t+\Delta t\mid\mathbf{x}_{i,k})-H_{k}(t\mid\mathbf{x}_{i,k})$, where $H_{k}(t\mid\mathbf{x}_{i,k})=\int_{0}^{t}h_{k}(u\mid\mathbf{x}_{i,k})du$ is the transition-specific cumulative hazard. Thus, it suffices to estimate $h_{k}(t\mid\mathbf{x}_{i,k})$ (Equation (1)) in order to compute the transition probability matrix. This approach permits flexible specification of covariate-dependent hazards, thereby inducing transition probabilities that vary with the underlying covariate structure [10]. By further including history-dependent information – such as state-entry times or multiple time scales – into the estimation of $h_{k}(t\mid\mathbf{x}_{i,k})$, Equation (2) can also handle non-Markov scenarios [11].
Traditional approaches to multi-state modeling, such as the non-parametric Aalen-Johansen (AJ) estimator [8] or cause-specific Cox Proportional Hazards (PH) models [12], have limited support for handling non-Markov scenarios and non-linear effect modeling. Furthermore, the application of multi-state models in the context of multi-stage disease histories comes with additional statistical challenges:
@@
    Independence between (history-dependent) state-entry times and events may be violated [13, 14, 15].
@@
    There exist multiple time scales, which are not all jointly identifiable in subpopulations due to the age-period-cohort problem [16].
@@
    Negative correlation between otherwise uncorrelated risk factors may be induced within subsets of diseased individuals, causing index event bias [17].
@@
    Disease histories derived from longitudinal data tend to be interval-censored as traits are only assessed at (ir-)regular follow-up times [18].
Here, we propose to use a novel multi-state modeling framework based on piecewise exponential additive models (PAMs) [19, 10]: This framework allows for seamless incorporation of history-dependent information and (time-varying) covariates as well as flexible estimation of their (time-varying) effects, for instance via penalized splines.
@@ 2.1 Dependent left-truncation
The multi-state process induces left-truncation because subjects enter the risk set for transition $k$ at different time points, i.e., at subject-specific state-entry times. An estimator for the transition-specific hazards $h_{k}(t\mid\mathbf{x}_{i,k})$ (Equation (1)) therefore needs to be able to take this into account. In general, left-truncation can be accounted for by non- and semi-parametric methods, provided that state-entry times (which represent the left-truncation times) can be assumed to be independent of event times. Such methods include the AJ estimator and the Cox PH model, which adjust the risk sets accordingly [20, 21].
However, state-entry times may not necessarily be independent of event times [13, 14, 15]: For instance, subjects with disease onset at a young age are possibly at higher risk of disease progression than subjects with late disease onset, implying dependence between truncation times and event occurrence. This dependence might not be linear on the log-hazard scale.
@@ 2.2 Choice of time scales
Hazards for all transitions of a multi-state model are commonly modeled on a single time scale, typically the natural time scale for transitions out of the initial state [16]. We call such a model with transition-specific stratification of the baseline hazard a stratified single time scale model. Risk sets for all transitions are defined along that single time scale $t$. Since the popular AJ estimator and Cox PH model can, by design, only incorporate a single time scale, the use of the stratified single time scale approach is further reinforced.
@@
In general registry data, entry into the dataset is random and not linked to the start of disease or intervention. Then, the baseline time scale is typically chronological age [22, 23] and age at disease onset is the time-to-event for the transition from ”healthy” into ”diseased”. For transitions into higher disease states or for clinical studies in patients (disease progression), the time since entry into the intermediate state is the relevant time scale [16, 24]. In clinical studies, time to recovery or relapse can be alternative time scales. This multiplicity of relevant time scales is illustrated for CKD onset and progression in Figure 2: For example, the hazard of progressing to Severe CKD for subject B at 57 years of age might depend not only on age, but also on time since CKD onset (i.e., 7 years). As a consequence, it may be of interest to allow transition intensities of a multi-state model to vary along multiple time scales; in the CKD example, time scales would be chronological age, time since CKD onset, and time since CKD progression. This way, baseline hazards can be captured by (a combination of) distinct time scales which commence at the time of entry into the different states [16, 25]. We call this the multiple time scales approach.
@@
Figure 2: Examples of disease histories regarding CKD. Based on the state diagram in Figure 1, this figure illustrates CKD onset and progression histories of three hypothetical subjects. Commencing at study entry (Healthy), the disease histories terminate with right-censoring or in the absorbing states ESKD or Death. Horizontal arrows represent time scales. The precise definitions of time scales and state-entry times, as well as the practical implementation via transition-specific helper variables, are described in Section 3.
@@
It is possible to jointly estimate the effects of multiple time scales and state-entry times by fitting one model on the entirety of disease trajectories, thus circumventing the identifiability problem of age-period-cohort models [16]. However, this requires fully parametric models with respect to parameter estimation, such as the piecewise exponential model [26] or discrete-time models [27]. The specific choice of (multiple) time scales can then be considered as an empirical – and thus, dataset-specific – question [16].
@@ 2.3 Index event bias
Diseases are typically multi-factorial, where risk factors can depend on each other [28]. Even if risk factors are independent in the healthy population, a negative correlation may be induced between them if analyses are restricted to subjects who have previously experienced a so-called index event whose occurrence is influenced by these risk factors. As an example, consider the index event CKD onset and two independent risk factors: the genetic variant rs77924615 from the UMOD locus (known to be associated with kidney function decline [29, 24]) and diabetes. When studying risk factor associations with CKD progression and thus restricting to CKD patients, the genetic variant will likely be associated with absence of diabetes (and vice versa) because people with the genetic variant are less likely to require an additional risk factor in order to be in the CKD subgroup.
As a consequence, failing to include other (confounding) risk factors into an analysis restricted to such a subgroup causes biased estimation of the effect of the risk factor of interest, in the form of attenuation or even effect reversal. This particular form of selection bias is referred to as index event bias (IEB) [17].
Conditioning on index events is inherent to multi-state models, due to their conditional view: For transitions out of all non-initial states $\tilde{\ell}\neq 0$, we condition on having previously transitioned into $\tilde{\ell}$. As a consequence, we expect all risk factors to be affected by IEB for progression analysis. In addition, as omitted variable bias (OVB) is known to affect time-to-event approaches by biasing effect size estimates towards zero [30], we additionally expect OVB-driven attenuation for all states.
@@ 2.4 Interval-censoring
Longitudinal datasets are routinely obtained either from cohort studies with prespecified examination times or from registry data with irregular and subject-specific visit numbers and times [31, 32, 33, 34, 35]. In both cases, we have a natural interval structure for each subject, implying that event times from longitudinal data are typically interval-censored [18].
The degree of interval-censoring (IC) inherently depends on interval lengths. For registry data, interval lengths tend to be smaller for less healthy individuals, who in turn are more prone to have events (”informative observation”) [36, 37]. This, as well as the fact that the common endpoint Death (cf. Figure 1) is usually recorded precisely [38], limits the potential extent of IC.

### v2

# Multi-state Models For Disease Histories Based On Longitudinal Data

@@ Abstract
Multi-stage disease histories derived from longitudinal data are becoming increasingly available as registry data and biobanks expand. Multi-state models are suitable to investigate transitions between different disease stages in presence of competing risks. In this context, however, their estimation is complicated by dependent left-truncation, multiple time scales, index event bias, and interval-censoring. In this work, we investigate the extension of piecewise exponential additive models (PAMs) to this setting and their applicability given the above challenges. In simulation studies we show that PAMs can handle dependent left-truncation and accommodate multiple time scales. Compared to a stratified single time scale model, a multiple time scales model is found to be less robust to the data generating process. We also quantify the extent of index event bias in multiple settings, demonstrating its dependence on the completeness of covariate adjustment. In general, PAMs recover baseline and fixed effects well in most settings, except for baseline hazards in interval-censored data. Finally, we apply our framework to estimate multi-state transition hazards and probabilities of chronic kidney disease (CKD) onset and progression in a UK Biobank dataset (n=$142,667$). We observe CKD progression risk to be highest for individuals with early CKD onset and to further increase over age. In addition, the well-known genetic variant rs77924615 in the UMOD locus is found to be associated with CKD onset hazards, but not with risk of further CKD progression.
@@ 1 Introduction
As the adoption of registry data expands [Slawomirski and others, 2023; Rau et al., 2024] and biobanks gain importance [Sudlow et al., 2015; of Us Research Program Investigators, 2019; Greiser et al., 2023], multi-stage disease histories are increasingly becoming available for use in both clinical routine and epidemiological research. Examples include Chronic Kidney Disease (CKD), where disease stages are defined by thresholds of serum creatinine-based estimated glomerular filtration rate (eGFR) measurements [Group, 2013]; diabetes, defined based on fasting plasma glucose levels [Committee, 2025]; or age-related macular degeneration (AMD), characterized by the accumulation of subretinal drusenoid deposits (early/intermediate AMD) and cell atrophy (late AMD).
In this context, it is usually of interest to study transitions and risk factors conditional on current disease stage: For example, given onset of Mild CKD at 50 years of age and progression to Severe CKD 10 years later, what is the probability of developing end-stage kidney disease (ESKD) within the next $t$ years? And is a genetic variant a risk factor for CKD onset, for CKD progression, or for both? It is appealing to employ multi-state models [cf. Andersen et al., 2012; Cook and Lawless, 2018] to derive transition probabilities and risk factor associations of multi-stage disease histories, such as CKD stages (Figure 1).
@@
Multi-stage disease histories can be formalized as a stochastic process $(X_{t})_{t\geq 0}$, where $X_{t}$ denotes the occupied state at time $t$ from a finite state space $\mathcal{S}:=\{0,1,\ldots,K\}$. As the process evolves over time, it induces a filtration $\mathcal{F}_{t}:=\sigma(X_{s}:0\leq s\leq t,\mathbf{x})$, representing the information available up to time $t$. This filtration comprises the observed trajectory of the process on $[0,t)$ as well as covariates $\mathbf{x}$, which may be time-dependent but are assumed to be observed prior to time $t$.
Within this framework [cf. Andersen et al., 2012; Beyersmann et al., 2011; Andersen and Keiding, 2002; Putter et al., 2007], for states $o,\ell\in\mathcal{S}$ with $o\neq\ell$ and times $s,t\in[0,\tau)$ such that $s<t$, multi-stage disease histories are characterized by transition probabilities
$p_{o\ell}(s,t)=\mathbb{P}(X_{t}=\ell\mid X_{s}=o,\mathcal{F}_{s}),$ (1)
or, equivalently, by transition intensities
$\alpha_{o\ell}(t)=\lim_{\Delta t\to 0}\frac{1}{\Delta t}p_{o\ell}(t,t+\Delta t).$ (2)
Assuming $\Delta t$ to be infinitesimally small and defining the probability to stay in state $o$ as $\alpha_{oo}(t)=-\sum_{\ell\neq o}\alpha_{o\ell}(t)$, the empirical transition probability matrix $\mathbf{P}(s,t)=\left(p_{o\ell}(s,t)\right)_{o,\ell\in\mathcal{S}}$ follows from the cumulative transition intensity $\text{d}\boldsymbol{\Lambda}=\left(\alpha_{o\ell}(t)\Delta t\right)_{o,\ell\in\mathcal{S}}$, and can be calculated via the product integral over time interval $[s,t)$ [Aalen and Johansen, 1978; Beyersmann et al., 2011], i.e.,
@@
Oftentimes, $(X_{t})_{t\geq 0}$ is assumed to be Markov, i.e., $\mathbb{P}(X_{t}=\ell\mid X_{s}=o,\mathcal{F}_{s})=\mathbb{P}(X_{t}=\ell\mid X_{s}=o)$; in other words, the transition probability is assumed to solely depend on the past via the current time $s$ and the occupied state. Traditional approaches to multi-state modeling, such as the non-parametric Aalen-Johansen (AJ) estimator [Aalen and Johansen, 1978] or cause-specific Cox Proportional Hazards (PH) models [Cox, 1972], have limited support for handling non-Markov scenarios and non-linear effect modeling. Within a Markov framework, recent Bayesian and continuous-time approaches have advanced the modeling of time-inhomogeneous disease progression [Williams et al., 2020; Luo et al., 2021; Kendall et al., 2025]. However, these methods remain fundamentally limited by the Markov assumption, which precludes the incorporation of subject-level history-dependent information. Alternative semi-Markov approaches, such as the Bayesian accelerated failure time (AFT) model [Klausch et al., 2023], have been developed to handle interval-censored data while relaxing restrictive Markov assumptions. However, these methods typically rely on parametric distribution families for transition times, which can entail severe bias in baseline hazard estimation if the underlying distribution is misspecified.
Furthermore, the application of multi-state models in the context of multi-stage disease histories comes with additional statistical challenges:
@@
    Independence between (history-dependent) state-entry times and events may be violated [Mackenzie, 2012; Cheng and Wang, 2015; Wang et al., 2024].
@@
    There exist multiple time scales, which are not all jointly identifiable in subpopulations due to the age-period-cohort problem [Iacobelli and Carstensen, 2013].
@@
    Negative correlation between otherwise uncorrelated risk factors may be induced within subsets of diseased individuals, causing index event bias [Dahabreh and Kent, 2011].
@@
    Disease histories derived from longitudinal data tend to be interval-censored as traits are only assessed at (ir-)regular follow-up times [Zhang and Sun, 2010].
Here, we propose to use a novel multi-state modeling framework based on piecewise exponential additive models (PAMs) [Bender and Scheipl, 2018; Piller et al., 2024]: Within this framework, multi-state data encodes transitions via tuples $(y^{entry}_{i,k},y^{exit}_{i,k},d_{i,k},\mathbf{x}_{i,k})$, where $i\in\{1,\ldots,n\}$ is the subject index, $k$ refers to the transition from state $o$ to $\ell$ (also noted as $o\rightarrow\ell$), $y^{entry}_{i,k}$ and $y^{exit}_{i,k}$ are risk set entry and exit times, due to either transitioning or censoring, $d_{i,k}$ is the transition-specific status indicator, and $\mathbf{x}_{i,k}$ is the subject- and transition-specific covariate vector [Piller et al., 2025]. Transition intensities $\alpha_{o\ell}(t)$ are then computed through transition-specific hazards $h_{k}(t\mid\mathbf{x}_{i,k})$ (see Equations (4), (5), and (7)). This permits flexible specification of covariate-dependent hazards, thereby inducing transition probabilities that vary with the underlying covariate structure [Piller et al., 2024]. By further including history-dependent information – such as state-entry times or multiple time scales – into the estimation of $h_{k}(t\mid\mathbf{x}_{i,k})$ through covariate adjustment, this PAM-based framework can explicitly handle non-Markov scenarios. Moreover, baseline hazards are estimated much more flexibly, without relying on distributional assumptions, yet still in a fully parametric fashion. Finally, the proposed framework allows for seamless incorporation of time-varying covariates as well as flexible estimation of their (time-varying) effects, for instance via penalized splines.
@@ 2.1 Dependent left-truncation
The multi-state process induces left-truncation because subjects enter the risk set for transition $k$ at different time points, i.e., at subject-specific state-entry times. For the above CKD example, a subject with CKD onset at age 50 years is only at risk of CKD progression after CKD onset and is thus left-truncated for CKD progression prior to the age of 50 (see subject B in Figure 2). An estimator for the transition-specific hazards $h_{k}(t\mid\mathbf{x}_{i,k})$ therefore needs to be able to take this into account. In general, left-truncation can be accounted for by non- and semi-parametric methods, provided that state-entry times (which represent the left-truncation times) can be assumed to be independent of event times. Such methods include the AJ estimator and the Cox PH model, which adjust the risk sets accordingly [Andersen et al., 2012; Zhang et al., 2009].
However, state-entry times may not necessarily be independent of event times [Mackenzie, 2012; Cheng and Wang, 2015; Wang et al., 2024]: For instance, subjects with disease onset at a young age are possibly at higher risk of disease progression than subjects with late disease onset, implying dependence between truncation times and event occurrence. This dependence might not be linear on the log-hazard scale.
@@ 2.2 Choice of time scales
Hazards for all transitions of a multi-state model are commonly modeled on a single time scale, typically the natural time scale for transitions out of the initial state [Iacobelli and Carstensen, 2013]. We call such a model with transition-specific stratification of the baseline hazard a stratified single time scale model. Risk sets for all transitions are defined along that single time scale $t$. Since the popular AJ estimator and Cox PH model can, by design, only incorporate a single time scale, the use of the stratified single time scale approach is further reinforced.
@@
In general registry data, entry into the dataset is random and not linked to the start of disease or intervention. Then, the baseline time scale is typically chronological age [Korn et al., 1997; Thiébaut and Bénichou, 2004] and age at disease onset is the time-to-event for the transition from ”healthy” into ”diseased”. For transitions into higher disease states or for clinical studies in patients (disease progression), the time since entry into the intermediate state is the relevant time scale [Iacobelli and Carstensen, 2013; Wiegrebe et al., 2024]. In clinical studies, time to recovery or relapse can be alternative time scales. This multiplicity of relevant time scales is illustrated for CKD onset and progression in Figure 2: For example, the hazard of progressing to Severe CKD for subject B at 57 years of age might depend not only on age, but also on time since CKD onset (i.e., 7 years). As a consequence, it may be of interest to allow transition intensities of a multi-state model to vary along multiple time scales; in the CKD example, the time scales are chronological age, time since CKD onset, and time since CKD progression. This way, baseline hazards can be captured by (a combination of) distinct time scales which commence at the time of entry into the different states [Iacobelli and Carstensen, 2013; Copenhagen, 2024]. We call this the multiple time scales approach.
@@
Figure 2: Examples of disease histories regarding CKD. Based on the state diagram in Figure 1, this figure illustrates CKD onset and progression histories of three hypothetical subjects. Commencing at study entry (Healthy), the disease histories terminate in the absorbing states ESKD or Death or cease to be observable following right-censoring. Horizontal arrows represent time scales. The precise definitions of time scales and state-entry times, as well as the practical implementation via transition-specific helper variables, are described in Section 3.
@@
It is possible to jointly estimate the effects of multiple time scales and state-entry times by fitting one model on the entirety of disease trajectories, thus circumventing the identifiability problem of age-period-cohort models [Iacobelli and Carstensen, 2013]. However, this requires fully parametric models with respect to parameter estimation, such as the piecewise exponential model [Friedman, 1982] or discrete-time models [Tutz et al., 2016]. The specific choice of (multiple) time scales can then be considered as an empirical – and thus, dataset-specific – question [Iacobelli and Carstensen, 2013].
@@ 2.3 Index event bias
Diseases are typically multi-factorial, where risk factors can depend on each other [Duarte et al., 2013]. Even if risk factors are independent in the healthy population, a negative correlation may be induced between them if analyses are restricted to subjects who have previously experienced a so-called index event whose occurrence is influenced by these risk factors. As an example, consider the index event CKD onset and two independent risk factors: the genetic variant rs77924615 from the UMOD locus (known to be associated with kidney function decline [Gorski et al., 2022; Wiegrebe et al., 2024]) and diabetes. When studying risk factor associations with CKD progression and thus restricting to CKD patients, the genetic variant will likely be associated with absence of diabetes (and vice versa) because people with the genetic variant are less likely to require an additional risk factor in order to be in the CKD subgroup.
As a consequence, failing to include other (confounding) risk factors into an analysis restricted to such a subgroup causes biased estimation of the effect of the risk factor of interest, in the form of attenuation or even effect reversal. This particular form of selection bias is referred to as index event bias (IEB) [Dahabreh and Kent, 2011].
Conditioning on index events is inherent to multi-state models, due to their conditional view: For transitions out of all non-initial states $\tilde{\ell}\neq 0$, we condition on having previously transitioned into $\tilde{\ell}$. As a consequence, we expect all risk factors to be affected by IEB for progression analysis. In addition, as omitted variable bias (OVB) is known to affect time-to-event approaches by biasing effect size estimates towards zero [Bretagnolle and Huber-Carol, 1988], we additionally expect OVB-driven attenuation for all states.
@@ 2.4 Interval-censoring
Longitudinal datasets are routinely obtained either from cohort studies with mostly prespecified examination times or from registry data with irregular and subject-specific visit numbers and times [Fitzmaurice et al., 2012; Diggle, 2002; Lokku et al., 2020; Hyun et al., 2017; Cheung et al., 2017]. In both cases, we have a natural interval structure for each subject, implying that event times from longitudinal data are typically interval-censored [Zhang and Sun, 2010].
The degree of interval-censoring (IC) inherently depends on interval lengths. For registry data, interval lengths tend to be smaller for less healthy individuals, who in turn are more prone to have events (”informative observation”) [Sperrin et al., 2017; Bible et al., 2024]. This, as well as the fact that the common endpoint Death (cf. Figure 1) is usually recorded precisely [”dual censoring”; Boruvka and Cook, 2016; Schober and Vetter, 2018], limits the potential extent of IC.